6x^4-19x@+15=0 solve for x?
Answer by mananth(16946)

6x^4-19x^2+15=0
6x^4-10x^2-9x^2+15=0
2x^2(3x^2-5)-3(3x^2-5)=0
(3x^2-5)(2x^2-3)=0
..
m.ananth@hotmail.ca